Supplemental Operations Report 22Apr2023

[title Supplemental Operations Report – April 22th]

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation.

Power system:
Solar: nominalSOC Last 24 hours: Max 100%; Min 61 %; Avg 78.4%.
VDC Last 24 hours: Max 59.12V; Min 46.35; Avg 51.61V.Generator run time: about 12am – 6am. Not used on April 15th.

Propane Reading Station Tank: 77%
Propane Reading Director Tank: 73%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 77%

Water (Static Tank) – 366 gallons
Water in GreenHab – 117 gallons
Water (Outpost tank) – 442 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: Yes, to Zubrin’s Head and back.
Hours: 190.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged. All corroded battery wires were received and replaced. Sojourner and Spirit need one additional short battery wire to be replaced. Parts are in transit.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. Picked up in Loa after repair, its alternator was replaced. To Hanksville for water runs and supplies pick up.

Crew Car used and why, where? Yes. From Grand Junction by crew. To Hanksville for water runs and supplies.

General notes and comments: Hab Car alternator was replaced.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Nominal

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Nominal. Hab heater was offline several times this week due to strong north winds. Current work is nominal after restart.

Summary of GreenHab Operations: Nominal.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation .

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Shade exterior painting completed. It will require a second coat in a month or so.
Old wood from the clothesline deck is in the process of being burned.
Director’s trailer:
1) Heater went offline after strong winds. The control board displayed a fault code (single red light flashing) – limit switch/airflow problem. Unplugging the control board from power resolved the problem.
2) Porch was leveled.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

GreenHab Report – April 21th

Crew 280 GreenHab Report 21Apr2023
GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 73.5F

Hours of supplemental light: N/A

Daily water usage for crops: 14 gallons

Daily water usage for research and/or other purposes: 1

Water in Blue Tank: 131 gallon

Time(s) of watering for crops: 09.30, 17:30 h

Changes to crops: N/A.

The general watering of plants and crops was done. The door was open the whole day with high temperatures and no wind. Neus Sabaté has taken 200gr of soil from Green Hab wheelbarrow to grow seeds for her Martian Batteries experiment at the Science Dome. One tomato plant has been moved where most of the others as during the night, the heater wind went directly to it.
The floor has been swept.

Harvest: 289 g cucumber (1 piece), 36 g micro greens, 163 g tomatoes, 1 g chives, 27g jazz mix, 1g salvia, and 1g thyme

Support/supplies needed

Sol Summary – April 21th

Crew 280 Sol Summary Report 21-04-2023

Sol: 5

Summary title: Together in Isolation

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: We kicked off the day with a nice breakfast followed by a 2-hour EVA to Roberts Rock Garden, where four crew members (Núria Jar, Neus Sabaté, Ariadna Farrés, and Cesca Cufí) explored the region with two main goals in mind: on the one hand, create a navigation map for future orientation challenges, and on the other, collect iron-rich rocks to create innovative Martian batteries.

At the end of the EVA, the team returned to the base, ready to refuel with a lunch prepared by Carla Conejo González (Executive Officer and Biologist) and Mariona Badenas-Agusti (Crew Commander & Astronomer). For this meal, the crew ate dehydrated scrambled eggs with a salad of greens and tomatoes from the GreenHab. In the afternoon, everyone worked on their research, with some members continuing to prepare their experimental setups.

Just before communicating with Mission Support, we enjoyed some homemade chocolate chip cookies. Tonight, Carla and Mariona have planned a “space” surprise for the team and everyone will watch a movie to welcome the start of the weekend. In a few days, Hypatia I will have spent a full week on Mars, conducting research, learning from each other, and adapting to a new lifestyle on the Red Planet.

Look ahead plan: The crew is looking forward to celebrating the day of St. George on Sunday. This Catalan celebration involves giving roses to friends, but given that roses do not exist on Mars, they will make them by hand with origami. In addition, the crew will organize a meditation workshop and an informal research symposium.

Anomalies in work: None.

Weather: Sunny. Low 35F/High 62F.

Crew physical status: Nominal.

EVA: We conducted an EVA to Roberts Rock Garden to collect iron oxide samples and plan a navigation GPS challenge.

Reports to be filed: Sol Summary, EVA request, EVA Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: Retrieval of trash in the rear airlock.

Operations Report – April 21th

Crew 280 Operations Report 21-04-2023

SOL: 5
Name of person filing report: Neus Sabaté (Mission Specialist) and Cesca Cufí (Crew Engineer))
Non-nominal systems:
Robotic observatory.
Rover Spirit.
Notes on non-nominal systems sections:
Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation. Corroded battery wire in Rover Spirit. Replacement pending (Sergii). Rover can only be used for short distance EVA.
Spirit rover used: yes
Hours: 217.0
Beginning charge: 100%
Ending charge: 65%
Currently charging: yes
Opportunity rover used: no
Hours: 121.1
Beginning charge: –
Ending charge: –
Currently charging: no
Curiosity rover used: no
Hours: 230.1
Beginning charge: 100%
Ending charge: 65%
Currently charging: yes
Perseverance rover used: yes
Hours: 261.5
Beginning charge: –
Ending charge: –
Currently charging: no
General notes on rovers: none
Summary of Hab operations:
Water use: 22 gal (from 6PM Sol4 to 6PM Sol5)
Water (static tank): 398 gal (at 6:00PM)
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 3, 6, 7 and 8.
Radios: EVA radios work nominally. Channel 1 com with the station was lost after passing by the road passage at Kissing Camel ridge, driving South. The EVA-Station com was maintained thanks to the crew’s own radio equipment (Midland HAM radios).
All radio and spacesuit batteries checked before EVA: ok.
Summary of GreenHab operations:
Water use: 14 gal for watering and 1 gal for experiments in Science Dome
Heater: off
Supplemental light: off
Harvest: 289 g cucumber (1 piece), 36 g micro greens, 163 g tomatoes, 1 g chives, 27g jazz mix, 1g salvia and 1g thyme
Neus Sabaté has taken 200gr of soil from Green Hab wheelbarrow to grow seeds for her Martian Batteries experiment at the Science Dome.
Summary of Science Dome operations:
Laia Ribas (Green Hab Officer): experiment.
Neus Sabaté (Mission Specialist): experiment.
Cesca Cufí (Crew Engineer): experiment
Dual split: off
Summary of RAM operations: CO monitor has been replaced.
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues:
Summary of other system’s issues: non
Questions, concerns and requests to Mission Support:
Today’s EVA has taken the team to Robert’s Garden Rock, which is relatively close to the base (approx. 1 mile) and the rover’s batteries discharged to 65%. We have observed that the batteries show a fast discharging rate at the beginning of the EVA and then keep discharging at a lower pace. Considering the Mission Support advice of not letting the rover’s batteries go down below 60% we would like to know if it is possible to plan a longer EVA that allows us to reach longer distances. Advice about this matter will be highly appreciated.
Spirit rover: replacement of the corroded wire
Kitchen: knobs replacement
Retrieval of trash in the rear airlock

Journalist Report – April 21th

Crew 280 Journalist Report 21-04-2023

Author: Núria Jar, Crew Journalist

Hypatia I is excited about the experiments all crew members are conducting. Collaboration and sharing ideas are essential for the success of their mission, especially in an isolated and confined environment, such as in the Mars Research Desert Station (MDRS).

Last days crew members were working together and sharing their first findings, challenges and solutions to ensure the success of their experiments. In this sense, the Hab table is an important location for the crew to come together and exchange these ideas, as it serves as a central hub for discussions, brainstorming and laughing out loud.

Recently, crew members became thrilled about the experiments conducted by the Mission Specialist, Neus Sabaté. Her research is focused on technologies that will enable life on Mars, particularly those related to energy. She is conducting an experiment to develop an iron-based battery using the compounds readily available in the surroundings of the MDRS. This was one of the purposes of the last extravehicular activity (EVA): to collect some rock samples, containing iron, to evaluate its use for Martian batteries.

Meanwhile, Sabaté is testing her system with commercial iron compounds to light up an LED system to germinate and grow edible greens at the Science Dome all night long. The rest of the crew members couldn’t wait to see the system. So, they put on headlamps to walk through the base corridors and make their way to the Science Dome to check out her system. As they gathered around the tiny lights, they felt together in isolation.

Also at the Science Dome, the Crew Engineer, Cesca Cufí-Prat, is building the Blob House, a single cell amoeba-like organism with learning capabilities. She wants to study the behavior and sporulation triggering of a single-cell organism under the influence of UV and infrared radiation, which is similar to the conditions found on the surface of Mars.

Next to her, the Crew Scientist, Ariadna Farrés Basiana, is developing a Martian GPS. During the EVA she collected some data to design navigation strategies for other Hypatia crew members. She spent much time taking notes.

In the kitchen, the Crew Commander, Mariona Badenas-Agustí, used some creative cooking methods to cook some delicious homemade chocolate chip cookies with dehydrated butter and eggs. They were a fantastic afternoon snack.

EVA Report – April 21th

Crew 280 EVA Report 21-04-2023

EVA #5

Author: Núria Jar (Crew Journalist) and Ariadna Farrés (HSO)

The EVA had four goals:
Evaluating the feasibility of navigation strategies developed by Ariadna Farrés (Scientist & HSO). Towards the end of the mission, as part of the navigation experiment, Ariadna would like to do an orientation exercise with the crew members; she needs to find locations to start planning this exercise.
Collecting some iron containing rock samples to evaluate it use for martian batteries by Neus Sabate (Mission Specialist); The project aims at demonstrating that it is possible to obtain Fe3+ ions from iron-containing rocks such as hematite, jarosite, limonite or argillite by immersing small pieces of rock in acid.
Documenting the rock sampling for outreach purposes, Núria Jar (Crew Journalist) filmed and took pictures of the crew while performing the EVA.
Creating a map of the sample collection area, Cesca Cufi (Crew Engineer) has experience in the field and will be creating a map of Roberts Rock Garden area, this has a double purpose: (1) help identify the areas that have already been sample for future expeditions; (2) help with the navigation and orientation exercise in future EVAs.

Start time: 10:14 AM

End time: 12:10 PM

Narrative: After performing 5 minutes of depressurization in the front airlock, the EVA team drove Spirit and Curiosity rovers to Roberts Rock Garden to perform the four main tasks: 1) evaluating the feasibility of navigation strategies; 2) collecting some iron containing rock; 3) documenting the rock sampling for outreach purposes; 4) and creating a map of the sample collection area.

After passing Kissing Camel Ridge the EVA crew lost communication with the base and were only able to communicate with the main Hab team with Hypatia’s own Midland radios. With the latter, communication between the EVA team and base was continuous.

When reaching our destination, Roberts Rock Garden, the rovers were parked and the team organized themselves into two groups. The first group, Neus Sabaté (Mission Specialist) and Núria Jar (Crew Journalist), explored the area looking for iron-containing rocks; they walked around the area around 4.5 miles, but never farther than 1000 meters from the rovers. Núria Jar documented all the sampling for outreach purposes. Unfortunately they did not find any iron-containing rocks. The second group Ariadna Farres (HSO) and Cesca Cufí (Crew Engineer) stayed closer to where the rovers were parked, identifying landmarks and planning the navigation exercise for the crew. The two groups maintained constant visual and radio communications.

After 1 hour exploring the area, the EVA crew walked back to the rovers, and started to drive back to the MDRS. The two rovers made a quick stop halfway before the main Hab, where they thought there might be iron-containing rocks, unfortunately this was not the case. The EVA ended an hour earlier than what was planned and will have to find other places to sample in future EVAs. Before entering the Hab, we followed the 5-minutes pressurization protocol.

Destination: Roberts Rock Garden

Coordinates (use UTM WSG 84): 518000 E, 4248000 N

EVA participants: 4: Núria Jar (Crew Journalist) (EVA Leader), Neus Sabaté (Mission Specialist), Ariadna Farrés (HSO), Cesca Cufí (Crew Engineer).

Road(s) and routes per MDRS Map: Drive on Cow Dung Road 0110

Mode of travel: Drive on Cow Dung Road 0110 to Roberts Rock Garden then walking around no farther than 1000 meters for sampling.

Copyright © The Mars Society. All rights reserved. | Main Site